OpenStudy (anonymous):

What is the solution to the equation log4 x + log 4 2 =3 i did x*2 = 3 2x = 3 x= 3/2 but my options are x=1/2 x=2 x=32 x=128

OpenStudy (phi):

\[ log_4(x) + log_4(2)= 3\] add logs is the same as multiplying and then taking the log \[ log_4(2x) = 3\] now make each side the exponent of 4 \[ 4^{log_4(2x)} = 4^3 \] the ugly stuff on the left turns into 2x (you should learn that this is how you "undo" a log) now solve for x \[ 2x= 4^3\] \[ x= \frac{1}{2}\cdot 64 \]

OpenStudy (anonymous):

32 :-)